OpenStudy (anonymous):

Write a compound inequality to represent all of the numbers between -4 and 6.

OpenStudy (anonymous):

Where are you stuck?

OpenStudy (anonymous):

- I'm not sure, this just doesn't make any sense to me ! I need help.

OpenStudy (anonymous):

-4<x<6 , in case of that problem you just need to put the lowest number in the beginning and always use the "<" and x and the end number, if you have a skip in domain or range you can put the union sign and add the next combination

OpenStudy (anonymous):

Let's say I want to write a compound inequality to represent the numbers from 0 to 10. It would look something like this: \[0 \le x \le 10\] Basically the x stand for any number. \[0 \le x\] means that 0 is below or equal to x, and \[x \le 10\] means that x is below or equal to 10, which means x must be higher than or equal to 0 AND lower than or equal to 10. Note: We *could* write this backwards like this: \[10 \ge x \ge 0\] but you'll almost never see it written that way...compound inequalities are pretty much always written with < signs, not > signs. Now, try writing the one that they want you to write.

OpenStudy (anonymous):

- K, thanks !